### v1

# arXiv:2603.00917v1[cs.CL] 01 Mar 2026

Abstract
Background: Artificial intelligence is increasingly deployed in healthcare workflows, and small open-source language models are gaining attention as viable tools for low-resource settings where cloud infrastructure is unavailable. Despite their growing accessibility, the reliability of these models — particularly the stability of their outputs under different phrasings of the same clinical question — remains poorly understood.
Objective: This study systematically evaluates prompt sensitivity and answer consistency in small open-source language models on clinical question answering benchmarks, with implications for low-resource healthcare deployment.
Methods: We evaluated five open-source models — Gemma 2 (2B), Phi-3 Mini (3.8B), Llama 3.2 (3B), Mistral 7B, and Meditron-7B (domain-pretrained, not instruction-tuned) — across three clinical question answering datasets (MedQA, MedMCQA, and PubMedQA), sampling 200 questions from each. Each question was presented in five prompt styles (original, formal, simplified, roleplay, and direct). We measured a majority-based consistency score, overall and per-style accuracy, and instruction-following failure rate (UNKNOWN rate). All inference was conducted locally on consumer CPU hardware without domain-specific fine-tuning.
Results: Consistency and accuracy were largely independent across models and datasets. Gemma 2 achieved the highest consistency scores (0.845–0.888) but the lowest accuracy (33.0–43.5%), while Llama 3.2 demonstrated moderate consistency (0.774–0.807) alongside the highest accuracy (49.0–65.0%). Roleplay prompts consistently reduced accuracy across all models and datasets, with Phi-3 Mini showing the largest decline of 21.5 percentage points on MedQA. Instruction-following failure rates varied by model and were not determined by parameter count, with Phi-3 Mini exhibiting the highest UNKNOWN rate at 10.5% on MedQA. Meditron-7B, a domain-pretrained model without instruction tuning, exhibited near-complete instruction-following failure on PubMedQA (99.0% UNKNOWN rate), demonstrating that domain knowledge alone is insufficient for structured clinical question answering.
Conclusions: High consistency does not imply correctness in small clinical language models — models can be reliably wrong, representing a particularly dangerous failure mode in clinical decision support. Roleplay prompt styles should be avoided in healthcare AI applications. Among models evaluated, Llama 3.2 demonstrated the strongest balance of accuracy and reliability for low-resource deployment. These findings highlight the necessity of multidimensional evaluation frameworks that assess consistency, accuracy, and instruction adherence jointly for safe clinical AI deployment.
Keywords: prompt sensitivity, answer consistency, small language models, clinical question answering, large language models, low-resource healthcare, instruction following, medical AI
1 Introduction
Artificial intelligence is increasingly being integrated into healthcare workflows, from clinical decision support to exam-style medical question answering. While much attention has focused on large, cloud-based models, smaller open-source language models are gaining importance because they can run locally on standard CPUs, making them viable for rural clinics, community hospitals, and low-resource health systems with limited infrastructure (Garg and others, 2024). However, as these smaller models become more accessible for real-world deployment, questions about their reliability become critical. In clinical settings, even small inaccuracies can have significant consequences, and the safety of AI systems depends not only on their overall performance but on how stable and predictable their outputs are under different conditions. Despite rapid adoption, the reliability of small, locally deployable models in healthcare remains insufficiently studied.
Most existing evaluations of clinical language models focus primarily on accuracy — whether a model selects the correct answer on a given benchmark (Bedi and others, 2025). However, accuracy alone does not capture how stable a model’s outputs are when the same clinical question is phrased differently. In safety-critical domains such as healthcare, this distinction matters: a model that gives different answers depending on how a question is worded cannot be considered reliable, regardless of its aggregate accuracy score.
To our knowledge, no prior study has systematically examined whether small, CPU-runnable models provide consistent answers under controlled prompt variation in clinical settings. The CLEVER framework (Kocaman and others, 2025) evaluates clinical LLM outputs through physician preference but does not assess response stability across prompt reformulations. Ngweta et al. (Ngweta and others, 2024) quantify prompt brittleness in general NLP tasks but do not examine clinical datasets or safety implications. Kim et al. (Kim and others, 2025) demonstrate that foundation models frequently produce confident but incorrect medical outputs, yet do not empirically measure variability under prompt changes. Recent surveys on small language models in healthcare (Garg and others, 2024) discuss efficiency and deployment tradeoffs without analysing answer consistency as a reliability dimension. This study addresses that gap directly.
In this work, we systematically evaluate prompt sensitivity in five open-source language models — Phi-3 Mini (3.8B), Llama 3.2 (3B), Gemma 2 (2B), Mistral 7B, and Meditron-7B (domain-pretrained) — across three established clinical question answering benchmarks (MedQA, MedMCQA, and PubMedQA) using five controlled prompt styles designed to simulate realistic query reformulation. We introduce a quantitative consistency score that measures how often a model produces the same answer across prompt variations for the same question, and analyse this alongside overall accuracy, per-style accuracy, and instruction-following failure rates. All experiments were conducted locally on consumer CPU hardware using unmodified base models without medical fine-tuning, reflecting realistic deployment conditions in low-resource healthcare settings.
We report four central findings. First, consistency and accuracy are independent metrics: models that produce highly stable answers across prompt variations are not necessarily more correct, with Gemma 2 demonstrating the clearest case of high consistency alongside low accuracy. Second, roleplay prompt styles systematically reduce performance across all models and datasets, indicating that persona-based framing negatively affects clinical question answering reliability. Third, instruction-following failure rates vary independently of model size, demonstrating that larger parameter counts do not guarantee safer or more dependable outputs under constrained deployment conditions. Fourth, Meditron-7B, a domain-pretrained model without instruction tuning, exhibited near-complete instruction-following failure despite encoding substantial medical knowledge, demonstrating that domain knowledge and instruction-following capability are orthogonal requirements for clinical deployment.
The remainder of this paper is structured as follows. Section 2 reviews related work on clinical LLM evaluation, prompt sensitivity, and small language models in healthcare. Section 3 describes our datasets, models, prompt variation design, inference setup, and evaluation metrics. Section 4 presents experimental results. Section 5 discusses clinical implications, limitations, and future directions. Section 6 concludes the paper.
2 Related Work
Most prior work evaluating large language models in healthcare has focused on accuracy, reasoning quality, and output safety. The CLEVER framework (Kocaman and others, 2025) evaluates clinical LLM outputs through blinded physician preference on tasks including summarisation and question answering, demonstrating that smaller domain-specific models can outperform larger general-purpose systems on clinically relevant criteria. Kim et al. (Kim and others, 2025) systematically examine medical hallucinations in foundation models, highlighting the dangers of confident but incorrect outputs and their potential to mislead clinical decision-making. While these studies provide critical insights into output quality and safety, neither investigates whether models produce consistent answers when the same clinical question is phrased differently — leaving prompt sensitivity largely unexplored in healthcare AI.
Growing interest in small, CPU-runnable language models has emerged in response to the practical constraints of healthcare deployment in low-resource settings. Garg et al. (Garg and others, 2024) survey small language models in healthcare, documenting their efficiency advantages and suitability for local deployment under limited infrastructure. The MedQA (Jin et al., 2021), MedMCQA (Pal et al., 2022), and PubMedQA (Jin et al., 2019) benchmarks provide standardised evaluation tasks across diverse clinical question formats, and have been widely adopted for assessing compact models. These works collectively advance the case for small model deployment in accessible healthcare AI, but focus primarily on task accuracy and computational efficiency rather than answer stability under prompt variation.
Prompt sensitivity — the tendency of model outputs to change in response to semantically equivalent but stylistically different inputs — has been studied in general NLP research. Ngweta et al. (Ngweta and others, 2024) demonstrate that minor changes in prompt formatting cause measurable performance fluctuations across multiple models and benchmarks, and propose mitigation strategies to reduce this brittleness. However, such investigations have not been extended to clinical question answering, and the safety implications of prompt sensitivity for healthcare AI deployment remain unexamined.
Taken together, these lines of research motivate the present study, which systematically evaluates prompt sensitivity and answer consistency in small open-source clinical language models — complementing prior work on accuracy, hallucination, and general prompt robustness with a focus on reliability under realistic deployment constraints.
3 Methods
3.1 Datasets
We evaluated model consistency and accuracy using three established clinical question answering benchmarks. MedQA (Jin et al., 2021) contains multiple-choice questions derived from the United States Medical Licensing Examination (USMLE), representing clinical knowledge required for medical licensure. MedMCQA (Pal et al., 2022) consists of multiple-choice questions from Indian medical entrance examinations (AIIMS/NEET), covering diverse medical specialties including pharmacology, anatomy, and pathology. Although the dataset includes subject labels, our evaluation used a randomly sampled subset of 200 questions without stratification by specialty. PubMedQA (Jin et al., 2019) presents biomedical research questions paired with PubMed abstracts, requiring yes, no, or maybe answers based on provided evidence rather than memorized knowledge. From each dataset we randomly sampled 200 questions using a fixed random seed (seed=42) to ensure reproducibility, yielding 600 questions total across three distinct clinical reasoning contexts.
3.2 Models
We evaluated five open-source small language models spanning a parameter range of 2B to 7B, including one domain-adapted medical pretraining model to examine the effect of model scale on prompt consistency. The selected models were Phi-3 Mini (Abdin and others, 2024) (3.8B parameters, Microsoft), Llama 3.2 (Grattafiori and others, 2024) (3B parameters, Meta), Gemma 2 (Gemma Team and others, 2024) (2B parameters, Google), and Mistral (Jiang and others, 2023) (7B parameters, Mistral AI). These models were selected because they are fully open-source, capable of running on consumer CPU hardware without GPU acceleration, and represent distinct model families — enabling cross-architecture comparison of consistency behavior. All inference was conducted locally using Ollama without any domain-specific fine-tuning, ensuring that observed consistency patterns reflect inherent model behavior rather than task-specific adaptation.
In addition to these instruction-tuned general models, we included Meditron-7B (Chen and others, 2023), a domain-adapted language model developed by EPFL. Meditron-7B is built on Llama-2 and continued-pretrained on a curated medical corpus comprising PubMed articles, biomedical abstracts, and internationally recognised clinical practice guidelines. Critically, Meditron-7B is a base pretraining model without instruction tuning, meaning it has not been trained to follow structured task instructions. We include it explicitly to examine whether domain-specific medical pretraining alone — in the absence of instruction tuning — is sufficient for structured clinical question answering, and to contrast its behaviour with that of general instruction-tuned models. This comparison addresses a practical question for clinical AI deployment: is domain knowledge alone sufficient, or is instruction-following capability equally essential?
3.3 Prompt Variation Design
To systematically evaluate how small language models respond to differently worded clinical prompts, we designed a prompt variation engine that generates five semantically equivalent but stylistically distinct prompt formulations for each question.
  • Original: The question exactly as it appears in the dataset, serving as the baseline condition.
  • Formal: Rephrased in clinical academic language to simulate how a healthcare professional might query a model in a structured setting.
  • Simplified: Written in plain everyday language, simulating how a non-expert or patient might pose the same question.
  • Roleplay: The model is instructed to respond as a practicing physician, testing whether persona-based prompting affects answer consistency.
  • Direct: Presents only the bare question and answer options with minimal framing, assessing model behavior without explicit instruction.
Each question across all three datasets was transformed into these five prompt styles, yielding 3,000 total prompts (600 questions $\times$ 5 styles). By keeping semantic content identical across variations, any differences in model output can be attributed to prompt sensitivity rather than question content.
3.4 Inference Setup
All models were served locally via Ollama and queried through its REST API. Inference was conducted with temperature set to 0 and a maximum token limit of 10 tokens per response. Setting temperature to 0 ensures fully deterministic outputs — the model produces identical responses for identical prompts — which is essential for isolating prompt sensitivity as the source of answer variation rather than stochastic sampling. Limiting output to 10 tokens prevents verbose responses and constrains the model to produce the single-letter (A, B, C, or D) or keyword (yes, no, or maybe) answer required by each dataset format.
Model responses were parsed using regular expression extraction. For multiple-choice datasets (MedQA and MedMCQA), the extractor identified the first standalone letter A, B, C, or D in the response. For PubMedQA, the extractor identified the first occurrence of yes, no, or maybe. Responses that did not contain a valid answer were categorized as UNKNOWN, representing instruction-following failures. All responses — including UNKNOWN outputs — were retained for consistency scoring and analysis.
3.5 Evaluation Metrics
We defined four complementary metrics to quantify model behavior under prompt variation.
Consistency Score. For each question, we identify the majority answer — the response selected most frequently across the five prompt styles — and compute the proportion of styles that agree with it. Formally, given responses $r_{1},r_{2},r_{3},r_{4},r_{5}$ for a question, the consistency score is:
$C=\frac{\sum_{i=1}^{5}\mathbf{1}[r_{i}=\hat{r}]}{5}$ (1)
where $\hat{r}$ is the majority answer and $\mathbf{1}[\cdot]$ is the indicator function. A score of 1.0 indicates perfect consistency across all prompt styles; a score of 0.2 indicates maximum inconsistency. For example, responses [A, B, B, B, A] yield a majority answer of B and a consistency score of 3/5 = 0.60.
Overall Accuracy. The majority answer for each question is compared against the dataset ground-truth label, reflecting task-level correctness while accounting for variability across prompt styles.
Per-Style Accuracy. Accuracy is computed separately for each prompt style (original, formal, simplified, roleplay, direct), enabling identification of which formulations the model handles most and least reliably.
UNKNOWN Rate. The proportion of responses that do not contain a valid answer option (A–D or yes/no/maybe), representing instruction-following failures that undermine clinical reliability.
Together these metrics provide a holistic assessment of model behavior: a model may be highly consistent but consistently wrong, accurate only under specific prompt styles, or prone to instruction-following failures. Evaluating any single metric in isolation would obscure these clinically relevant failure modes.
4 Results
4.1 Overall Consistency and Accuracy
[…]
Dataset Model Mean Fully Overall UNKNOWN Acc. Acc. Acc. Acc. Acc.
Cons. Cons. Acc. Rate Orig. Form. Simp. Role. Dir.
MedQA Phi-3 Mini 0.698 19.5 48.0 10.5 44.5 44.5 42.5 26.5 48.0
Llama 3.2 0.776 35.0 49.0 0.8 47.5 44.5 46.0 47.0 43.5
Gemma 2 0.888 63.5 40.0 2.1 39.0 39.0 37.0 35.0 35.0
Mistral 7B 0.800 41.0 45.0 4.7 45.5 41.5 43.0 34.0 50.5
[…]
MedMCQA Phi-3 Mini 0.730 27.0 53.0 4.6 52.0 49.5 48.5 37.5 52.0
Llama 3.2 0.774 38.5 55.5 1.5 56.5 50.5 51.5 48.0 53.0
Gemma 2 0.851 56.5 43.5 0.9 42.5 39.0 38.0 40.5 44.0
Mistral 7B 0.812 45.5 45.5 4.7 44.5 43.0 42.5 41.5 44.0
[…]
PubMedQA Phi-3 Mini 0.830 41.0 48.0 0.0 42.5 53.5 52.5 47.5 31.0
Llama 3.2 0.807 45.5 65.0 2.8 59.5 51.0 64.5 65.0 56.5
Gemma 2 0.845 50.5 33.0 1.3 28.0 41.5 47.5 27.0 30.5
Mistral 7B 0.825 49.5 42.5 7.2 39.5 39.5 45.5 34.5 45.5
[…]
Statistical significance of observed differences was assessed using the Wilcoxon signed-rank test for consistency scores and McNemar’s test for accuracy comparisons. Consistency differences between models were statistically significant on MedQA and MedMCQA (p<<0.001 for most pairwise comparisons) but not on PubMedQA, suggesting that dataset characteristics influence the degree of prompt sensitivity. Accuracy differences were generally non-significant on MedQA, while on PubMedQA, Llama 3.2 significantly outperformed both Phi-3 Mini and Gemma 2 (p<<0.001), confirming its superior clinical reasoning on biomedical evidence-based questions. Full significance test results are provided in Table LABEL:tab:stats.
[…]
Dataset M1 M2 Metric p Sig.
MedQA Llama 3.2 Phi-3 Cons. 0.0005 ***
Gemma 2 Phi-3 Cons. <<.0001 ***
Mistral Phi-3 Cons. <<.0001 ***
Llama 3.2 Gemma 2 Cons. <<.0001 ***
Meditron Phi-3 Cons. <<.0001 ***
Meditron Llama 3.2 Cons. 0.3342 ns
Llama 3.2 Phi-3 Acc. 0.9020 ns
Gemma 2 Phi-3 Acc. 0.0689 ns
Mistral Phi-3 Acc. 0.5383 ns
Llama 3.2 Gemma 2 Acc. 0.0512 ns
Meditron Phi-3 Acc. 0.0004 ***
Meditron Llama 3.2 Acc. 0.0001 ***
MedMCQA Llama 3.2 Phi-3 Cons. 0.0437 *
Gemma 2 Phi-3 Cons. <<.0001 ***
Mistral Phi-3 Cons. 0.0001 ***
Llama 3.2 Gemma 2 Cons. 0.0001 ***
Meditron Phi-3 Cons. 0.0007 ***
Meditron Llama 3.2 Cons. 0.2052 ns
Llama 3.2 Phi-3 Acc. 0.5962 ns
Gemma 2 Phi-3 Acc. 0.0327 *
Mistral Phi-3 Acc. 0.1060 ns
Llama 3.2 Gemma 2 Acc. 0.0075 **
Meditron Phi-3 Acc. 0.0002 ***
Meditron Llama 3.2 Acc. <<.0001 ***
PubMedQA Llama 3.2 Phi-3 Cons. 0.1055 ns
Gemma 2 Phi-3 Cons. 0.8497 ns
Mistral Phi-3 Cons. 0.6172 ns
Llama 3.2 Gemma 2 Cons. 0.1055 ns
[…]
Llama 3.2 Phi-3 Acc. <<.0001 ***
Gemma 2 Phi-3 Acc. 0.0002 ***
Mistral Phi-3 Acc. 0.1930 ns
Llama 3.2 Gemma 2 Acc. <<.0001 ***
Meditron Phi-3 Acc. <<.0001 ***
Meditron Llama 3.2 Acc. <<.0001 ***
Across all three datasets, consistency scores varied across models without a clear correlation with model size (Figure 1). Phi-3 Mini (3.8B) exhibited the lowest mean consistency scores (0.698–0.830), while Gemma 2 (2B) achieved the highest (0.845–0.888) despite being the smallest model tested. Mistral 7B demonstrated strong but slightly variable consistency (0.800–0.825), and Llama 3.2 showed intermediate consistency (0.774–0.807).
Accuracy patterns diverged substantially from consistency rankings (Figure 2). Llama 3.2 achieved the highest overall accuracy across datasets (49.0–65.0%), followed by Phi-3 Mini (48.0–53.0%) and Mistral 7B (42.5–45.5%). Gemma 2, despite exhibiting the highest consistency scores, achieved the lowest accuracy (33.0–43.5%). This inverse relationship between consistency and accuracy in Gemma 2 highlights a critical finding: high consistency does not imply high accuracy. Models can be reliably wrong — producing the same incorrect answer across all prompt variations — which represents a particularly dangerous failure mode in clinical decision support contexts.
Meditron-7B, included as a domain-adapted pretraining model without instruction tuning, exhibited markedly different behaviour from all instruction-tuned models. While it achieved moderate consistency on MedQA and MedMCQA (0.761 and 0.762 respectively), its UNKNOWN rates were substantially higher (22.8% and 22.4%), indicating frequent instruction-following failures. On PubMedQA, Meditron-7B produced near-complete failure — a consistency score of 0.010, 2.0% accuracy, and 99.0% UNKNOWN rate — suggesting the model is unable to produce structured yes/no/maybe responses without instruction tuning. These results indicate that domain-specific pretraining alone is insufficient for structured clinical question answering, and that instruction-following capability is a prerequisite for clinical deployment.
Figure 1: Mean consistency scores across models and datasets. Higher scores indicate greater agreement across prompt styles.
Figure 2: Overall accuracy across models and datasets, calculated using majority answer against ground truth labels.
4.2 Effect of Prompt Style on Accuracy
Prompt style demonstrated a consistent and measurable effect on model accuracy across all five models and three datasets (Figure 3). The Roleplay prompt style — in which models were instructed to respond as a practicing physician — consistently underperformed relative to all other styles. The most pronounced decline was observed in Phi-3 Mini, which achieved 48.0% accuracy under the Direct style but only 26.5% under Roleplay on MedQA, representing a drop of 21.5 percentage points. Similar reductions were observed across Llama 3.2, Gemma 2, and Mistral 7B, confirming that roleplay prompting is systematically detrimental rather than model-specific (Figure 4).
In contrast, the Direct and Original prompt styles produced the highest and most stable accuracy across models and datasets, suggesting that minimal framing is preferable for reliable clinical question answering. These findings carry direct implications for healthcare AI deployment: prompt formulations intended to simulate clinical expertise — such as persona-based instructions — may paradoxically reduce the reliability of small language model outputs in patient-facing applications.
Figure 3: Accuracy by prompt style across models and datasets. Roleplay consistently underperforms relative to other styles.
Figure 4: Roleplay prompt accuracy versus best performing non-roleplay style across all models and datasets, with annotated performance gaps.
4.3 Instruction-Following Failure Rate
The UNKNOWN rate — the proportion of responses that did not contain a valid answer option — varied substantially across models and datasets (Figure 5). Phi-3 Mini exhibited the highest UNKNOWN rate on MedQA at 10.5%, indicating that approximately one in ten queries failed to produce a usable response. Mistral 7B, despite being the largest model evaluated, also demonstrated non-negligible UNKNOWN rates across all three datasets (4.7%–7.2%), suggesting that model size alone does not guarantee reliable instruction following.
In contrast, Llama 3.2 and Gemma 2 achieved the lowest UNKNOWN rates across datasets (0.8%–2.8% and 0.9%–2.1% respectively), consistently producing responses in the expected format. These results suggest that instruction adherence is a model-specific characteristic independent of parameter count.
Meditron-7B exhibited the most severe instruction-following failures of all models evaluated, with UNKNOWN rates of 22.8% on MedQA, 22.4% on MedMCQA, and 99.0% on PubMedQA. The near-complete failure on PubMedQA — where 99 of 100 responses were invalid — is consistent with Meditron’s architecture as a base pretraining model without instruction tuning. Unlike the instruction-following failures observed in Phi-3 Mini, which appear to be prompt-style dependent, Meditron’s failures reflect a fundamental absence of structured response capability rather than prompt sensitivity.
From a clinical deployment perspective, high UNKNOWN rates represent a critical reliability concern. A model that frequently fails to produce a valid response cannot serve as a dependable decision support tool, as such failures interrupt clinical workflows, introduce uncertainty, and erode clinician trust in AI-assisted systems.
Figure 5: Instruction-following failure rates (UNKNOWN responses) across models and datasets. Phi-3 Mini shows the highest failure rate on MedQA at 10.5%.
4.4 Consistency Versus Accuracy Relationship
Analysis of the relationship between mean consistency scores and overall accuracy revealed no clear positive correlation across models and datasets (Figure 6). Gemma 2 exemplifies this disconnect most starkly: it achieved the highest consistency scores across all datasets (0.845–0.888) while simultaneously recording the lowest accuracy (33.0–43.5%), indicating that it consistently produces incorrect answers with high confidence. Conversely, Llama 3.2 demonstrated moderate consistency (0.774–0.807) yet achieved the highest accuracy across most datasets (49.0–65.0%), showing that a model can be more frequently correct even when its answers vary somewhat across prompt styles.
The distribution of consistency scores further illustrates model-level differences in reliability (Figure 7). Phi-3 Mini exhibited the widest spread of consistency scores on MedQA, indicating high variability in prompt sensitivity across questions. Gemma 2 showed a tighter, higher distribution — consistently stable but systematically inaccurate.
These findings demonstrate that consistency and accuracy are independent dimensions of model performance. Evaluating either metric in isolation is insufficient for clinical AI assessment. A model that is highly consistent but consistently wrong may be more dangerous than one that occasionally contradicts itself, as the former provides false confidence in systematically incorrect outputs. Safe clinical deployment therefore requires joint evaluation of both consistency and accuracy.
Figure 6: Scatter plot of mean consistency score versus overall accuracy across all models and datasets. No clear positive correlation is observed, indicating that consistency and accuracy are independent metrics.
Figure 7: Distribution of consistency scores across models and datasets. Phi-3 Mini shows the widest spread on MedQA; Gemma 2 shows a tight but high distribution reflecting systematic consistency without accuracy.
5 Discussion
5.1 Consistency and Accuracy are Independent Metrics
A central finding of this study is that consistency and accuracy represent independent dimensions of small language model performance in clinical question answering. While prior evaluations of clinical AI have predominantly focused on accuracy as the primary metric (Bedi and others, 2025), our results demonstrate that a model can achieve high consistency while remaining systematically inaccurate — a failure mode we term reliable incorrectness.
Gemma 2 illustrates this most clearly, achieving the highest mean consistency scores across all datasets (0.845–0.888) while recording the lowest overall accuracy (33.0–43.5%). This pattern suggests that the model has learned stable response tendencies that are nonetheless clinically wrong.
The danger of reliable incorrectness is as much psychological as technical. Consider a primary care physician using an AI assistant to support diagnostic reasoning. If the model returns the same answer regardless of how a question is phrased, this repetition creates an illusion of reliability. A clinician might reason: “it keeps suggesting pulmonary embolism — it must be confident.” That apparent stability feels reassuring. But if the correct diagnosis is pneumonia, the model’s consistency actively reinforces a false conclusion. Over time, clinicians may develop misplaced trust in systems that appear stable and decisive without recognising that the stability reflects systematic error rather than clinical knowledge. In practice, this could translate into repeated misdiagnoses, inappropriate investigations, or delayed treatment — not because the model behaves randomly, but because it is confidently and predictably incorrect.
This finding aligns with Kim et al. (Kim and others, 2025), who demonstrate that foundation models frequently produce hallucinations with high apparent confidence, and that clinicians consider such outputs particularly dangerous precisely because they are difficult to detect. Our results extend this concern to the domain of prompt sensitivity: a model that produces consistent but wrong answers across varied prompt formulations compounds this risk by appearing robust to rephrasing.
These observations underscore that accuracy and consistency must be evaluated jointly in any clinical AI assessment framework. Neither metric alone is sufficient to characterise the safety profile of a model intended for healthcare deployment.
5.2 Roleplay Prompts Consistently Underperform
Across all five models and three datasets, roleplay-style prompts consistently produced lower accuracy than direct or original question formats. The most pronounced decline was observed in Phi-3 Mini on MedQA, where roleplay accuracy dropped 21.5 percentage points below the best-performing style. Critically, this pattern was consistent across all model families tested, suggesting a systematic rather than model-specific phenomenon.
We propose that roleplay prompts introduce a form of task interference that is particularly detrimental for small language models. When a prompt begins with persona framing such as “You are a senior physician taking a licensing examination,” the model must simultaneously interpret and simulate a clinical identity while reasoning through a structured medical question. For smaller models with limited representational capacity, this dual requirement may dilute attention from the core reasoning task. Furthermore, persona-based prompts may activate patterns learned from conversational or narrative training data rather than the structured exam-style reasoning required for clinical QA benchmarks — effectively shifting the task from question answering to performance-based text generation.
This finding has direct practical implications. Prompt engineering guidelines for clinical AI systems often recommend persona framing to make outputs sound more authoritative or clinically appropriate (White and others, 2023). Our results suggest that for small open-source models, such framing may paradoxically reduce factual reliability. Developers building clinical decision support tools on small language models should favour minimal, direct prompt formulations over persona-based instructions.
This observation also extends findings from Ngweta et al. (Ngweta and others, 2024), who demonstrated that prompt format changes cause measurable performance variation in general NLP tasks. Our results show that in clinical settings, one specific format — roleplay — is systematically and substantially more harmful than others, representing a concrete and actionable finding for healthcare AI deployment.
5.3 Implications for Low-Resource Healthcare Deployment
A primary motivation for this study was the practical challenge of deploying AI in resource-constrained healthcare environments. Many rural clinics, community hospitals, and health systems in low-income settings lack access to GPU infrastructure, reliable high-speed internet, or cloud-based AI services. In these contexts, locally deployable open-source models in the 2B–7B parameter range represent a more realistic and equitable alternative to large proprietary systems. Our experimental design intentionally reflected these constraints: all models were evaluated on consumer CPU hardware without domain-specific fine-tuning, simulating realistic deployment conditions rather than idealised research environments.
Our findings provide concrete guidance for model selection in such settings. Although Gemma 2 demonstrated the highest consistency scores, its substantially lower accuracy suggests it may systematically mislead clinicians — a risk that is amplified in low-oversight environments where AI outputs may not be routinely verified. Mistral 7B, despite being the largest model evaluated, did not clearly outperform smaller alternatives and exhibited unexpected instruction-following failures, suggesting that larger parameter counts do not automatically justify greater hardware demands.
Llama 3.2 (3B) emerged as the most balanced candidate across our evaluation criteria, achieving the highest overall accuracy (49.0–65.0%), moderate consistency, and among the lowest instruction-following failure rates. For low-resource healthcare deployment scenarios where clinician oversight may be limited, a model offering strong accuracy and reliable instruction adherence is preferable to one that maximises consistency at the expense of correctness.
More broadly, our results suggest that deployment decisions for clinical AI systems should be guided by joint evaluation of accuracy, consistency, and instruction compliance rather than any single metric. In settings where errors carry direct patient safety implications, the cost of systematic incorrectness — even when delivered consistently — outweighs the benefit of apparent stability. This study provides a practical evaluation framework for low-resource healthcare AI deployment that can be reproduced without specialised infrastructure, making it accessible to researchers and practitioners in exactly the settings it is designed to serve.
5.4 Domain Knowledge Versus Instruction Following
The inclusion of Meditron-7B reveals an important distinction between two capabilities required for clinical AI deployment: domain knowledge and instruction adherence. Meditron-7B was pretrained on a curated medical corpus comprising PubMed articles and clinical guidelines (Chen and others, 2023), encoding substantial medical domain knowledge. However, without instruction tuning, it failed to produce structured responses in the required format across all three datasets, with UNKNOWN rates reaching 99.0% on PubMedQA.
This finding suggests that domain knowledge and task usability are orthogonal capabilities. A model can encode rich medical knowledge through continued pretraining yet remain clinically unusable for structured question answering without the ability to follow task-specific instructions. For practitioners deploying AI in clinical settings, this distinction has direct practical implications: selecting a model based on its training data alone — without verifying instruction-following capability — may result in a system that is knowledgeable but functionally unusable in structured workflows.
This also motivates a clear direction for future work: instruction-tuning Meditron or similar domain-pretrained models on structured clinical QA data, and evaluating whether the combination of domain knowledge and instruction following produces models that are both more accurate and more consistent than general instruction-tuned alternatives.
5.5 Limitations
Several limitations should be considered when interpreting these findings. First, we sampled 200 questions per dataset, which provides a computationally feasible and reproducible evaluation but may not fully capture the diversity of each benchmark. Second, the three datasets used — MedQA, MedMCQA, and PubMedQA — focus primarily on structured examination-style and research-based questions, which may not reflect the complexity of real-world clinical dialogue or unstructured patient interactions.
Third, we did not conduct human clinical evaluation to assess whether consistent but incorrect model outputs would meaningfully influence physician decision-making in practice. Such evaluation would strengthen the clinical validity of our findings but was beyond the scope of this study. Fourth, all models were evaluated without domain-specific fine-tuning, which may underestimate performance achievable through medical adaptation techniques such as instruction tuning or retrieval-augmented generation (Lewis and others, 2020).
Fifth, our evaluation was restricted to multiple-choice and yes/no/maybe answer formats, whereas real clinical environments frequently involve open-ended reasoning, differential diagnosis, and contextual patient data. Sixth, hardware constraints limited our evaluation to models in the 2B–7B parameter range running on consumer CPU hardware, precluding direct comparison with larger proprietary systems such as GPT-4 or Claude. Seventh, while our five prompt styles were designed to simulate realistic variation in clinical query formulation, they do not exhaust the full range of phrasings a clinician might use in practice. Finally, Meditron-7B was evaluated in its base pretraining form without instruction tuning, which limits direct comparison with instruction-tuned models and may underrepresent the potential of domain-adapted models when properly instruction-tuned for structured clinical tasks.
Together, these limitations suggest that our results should be interpreted as an initial systematic analysis of prompt sensitivity under constrained deployment conditions rather than a comprehensive evaluation of all clinical AI scenarios. Future work should address these gaps through larger samples, broader dataset coverage, and human clinical validation studies.
5.6 Future Work
Several directions emerge naturally from this study. First, future research should evaluate whether domain-specific fine-tuning of small models on medical corpora reduces prompt sensitivity and improves consistency. If fine-tuned models demonstrate stronger stability under prompt variation, this would suggest that training data alignment — rather than model architecture — is the primary driver of inconsistency.
In particular, instruction-tuning Meditron-7B or similar domain-pretrained models on structured clinical QA data represents a direct next step — testing whether combining domain knowledge with instruction-following capability produces models that are both more accurate and more consistent than general instruction-tuned alternatives evaluated in this study.
Second, integrating retrieval-augmented generation (RAG) pipelines with small clinical models represents a promising direction. RAG-augmented systems ground responses in retrieved evidence, which may reduce both factual hallucinations and prompt-induced answer variation. Testing whether external knowledge retrieval mitigates the consistency-accuracy tradeoff observed in this study would be a natural extension (Lewis and others, 2020).
Third, evaluating larger open-source and proprietary models — when hardware resources permit — would clarify whether the independence between consistency and accuracy observed here persists at greater parameter scales, or whether it is specific to sub-7B models.
Beyond technical extensions, human-centered validation is essential. A controlled clinical study involving practicing physicians could assess how consistent but incorrect model outputs influence clinical reasoning, trust calibration, and decision-making in realistic workflows. Such a study would strengthen the ecological validity of the consistency metric introduced here.
Translationally, our findings support the development of a lightweight clinical decision support prototype for general practitioners in low-resource settings, built around the model demonstrating the strongest balance of accuracy and instruction adherence, with prompt engineering guardrails informed by our sensitivity analysis. Finally, extending this evaluation framework to multilingual datasets and non-English clinical benchmarks is critical, as many low-resource healthcare environments operate outside English-dominant settings. Together, these directions would advance this work from controlled benchmarking toward safe, deployable clinical AI systems grounded in both technical rigour and real-world constraints.
6 Conclusion
In this study, we systematically evaluated prompt sensitivity in five open-source language models (2B–7B parameters), including four general instruction-tuned models and one domain-adapted pretraining model, across three clinical question answering datasets using five controlled prompt style variations, yielding 3,000 total evaluation instances.
We report four principal findings. First, consistency and accuracy are independent metrics: models that achieve high consistency are not necessarily more accurate, with Gemma 2 demonstrating the most pronounced case of reliable incorrectness — achieving the highest consistency scores while recording the lowest accuracy across all datasets. Second, roleplay-style prompts consistently reduced accuracy across all models and datasets, suggesting that persona-based framing introduces systematic instability in clinical reasoning tasks and should be avoided in healthcare AI prompt design. Third, instruction-following failures varied substantially across models and were not determined by parameter count alone, indicating that model size is an insufficient proxy for deployment reliability under hardware constraints. Fourth, Meditron-7B, a domain-adapted model without instruction tuning, exhibited near-complete failure on structured clinical QA tasks despite encoding substantial medical domain knowledge — demonstrating that domain knowledge and instruction-following capability are orthogonal requirements for clinical AI deployment.
Taken together, these findings demonstrate that consistency alone cannot serve as a proxy for clinical reliability. Models that appear stable under prompt variation may still pose significant patient safety risks if their answers are systematically incorrect. Safe deployment of AI in clinical settings — particularly in low-resource environments where human oversight may be limited — requires joint evaluation of accuracy, consistency, and instruction adherence.
We call on the research community to move beyond accuracy-only evaluation frameworks for clinical AI and adopt multidimensional assessment approaches that reflect the reliability requirements of real-world healthcare deployment. The evaluation framework, datasets, and code introduced in this study are publicly available at https://github.com/shravani-01/clinical-llm-eval to support reproducibility and future research in this direction.