### v3

# Prompt Sensitivity and Answer Consistency in Small Open-Source Language Models for Clinical Question Answering: An Empirical Evaluation for Low-Resource Healthcare Deployment

Abstract
Background: Artificial intelligence is increasingly deployed in healthcare workflows, and small open-source language models are gaining attention as viable tools for low-resource settings where cloud infrastructure is unavailable. Despite their growing accessibility, the reliability of these models, particularly the stability of their outputs under different phrasings of the same clinical question, remains poorly understood.
Objective: This study systematically evaluates prompt sensitivity and answer consistency in small open-source language models on clinical question answering benchmarks, with implications for low-resource healthcare deployment.
Methods: Five open-source language models spanning distinct architectural and training paradigms were evaluated across three clinical question answering datasets (MedQA, MedMCQA, PubMedQA) using five controlled prompt style variations, yielding 15,000 total inference calls conducted locally on consumer CPU hardware without fine-tuning. Consistency scores, accuracy, and instruction-following failure rates were measured and interpreted in the context of each model’s architectural design.
Results: Consistency and accuracy were largely independent across models and datasets. Gemma 2 achieved the highest consistency scores (0.845–0.888) but the lowest accuracy (33.0–43.5%), producing perfectly consistent yet incorrect answers on 77 of 200 MedQA questions (38.5%), a failure mode termed reliable incorrectness. Llama 3.2 demonstrated moderate consistency (0.774–0.807) alongside the highest accuracy (49.0–65.0%). Roleplay prompts consistently reduced accuracy across all models and datasets, with Phi-3 Mini showing the largest decline of 21.5 percentage points on MedQA. Instruction-following failure rates varied by model and were not determined by parameter count, with Phi-3 Mini exhibiting the highest UNKNOWN rate at 10.5% on MedQA. Meditron-7B, a domain-pretrained model without instruction tuning, exhibited near-complete instruction-following failure on PubMedQA (99.0% UNKNOWN rate), demonstrating that domain knowledge alone is insufficient for structured clinical question answering.
Conclusions: High consistency does not imply correctness in small clinical language models; models can be reliably incorrect, representing a potentially dangerous failure mode in clinical decision support. Roleplay prompt styles should be avoided in healthcare AI applications. Among the models evaluated, Llama 3.2 demonstrated the strongest balance of accuracy and reliability for low-resource deployment. These findings highlight the necessity of multidimensional evaluation frameworks that assess consistency, accuracy, and instruction adherence jointly for safe clinical AI deployment.
Keywords: prompt sensitivity, answer consistency, small language models, clinical question answering, large language models, low-resource healthcare, instruction following, medical AI
1 Introduction
Artificial intelligence is increasingly being integrated into healthcare workflows, including clinical decision support, medical documentation, and exam-style clinical question answering. While much attention has focused on large cloud-based models, smaller open-source language models are gaining importance because they can run locally on standard CPUs, making them viable for rural clinics, community hospitals, and low-resource health systems with limited infrastructure (Garg and others, 2024). As these smaller models become more accessible for real-world deployment, questions about their reliability become critical. In clinical settings, even small inaccuracies can have significant consequences, and the safety of AI systems depends not only on their overall performance but also on how stable and predictable their outputs are under different conditions. Despite increasing adoption, the reliability of small, locally deployable models in healthcare remains insufficiently studied.
Most existing evaluations of clinical language models focus primarily on accuracy, that is, whether a model selects the correct answer on a given benchmark (Bedi and others, 2025). However, accuracy alone does not capture how stable a model’s outputs are when the same clinical question is phrased differently. In safety-critical domains such as healthcare, this distinction is important. A model that gives different answers depending on how a question is worded cannot be considered reliable, regardless of its aggregate accuracy score.
Prior work has not systematically evaluated whether small, CPU-runnable language models provide consistent answers under controlled prompt variation in clinical contexts. The CLEVER framework (Kocaman and others, 2025) evaluates clinical LLM outputs through physician preference but does not assess response stability across prompt reformulations. Ngweta et al. (Ngweta and others, 2024) quantify prompt brittleness in general NLP tasks but do not examine clinical datasets or safety implications. Kim et al. (Kim and others, 2025) demonstrate that foundation models frequently produce confident but incorrect medical outputs, yet do not empirically measure variability under prompt changes. Surveys on small language models in healthcare (Garg and others, 2024) discuss efficiency and deployment tradeoffs but do not analyze answer consistency as a reliability dimension. The present study addresses this gap.
This study addresses three research questions:
RQ1: Do small open-source language models produce consistent answers when semantically equivalent clinical questions are phrased differently?
RQ2: Is answer consistency correlated with clinical question answering accuracy?
RQ3: How do different prompt styles influence the reliability of model outputs?
To investigate these questions, this work evaluates five open-source language models—Phi-3 Mini (3.8B), Llama 3.2 (3B), Gemma 2 (2B), Mistral 7B, and Meditron-7B (domain-pretrained)— across three established clinical question answering benchmarks (MedQA, MedMCQA, and PubMedQA). Each question is evaluated under five controlled prompt styles designed to simulate realistic query reformulation. A quantitative consistency score is introduced to measure how often a model produces the same answer across prompt variations for the same question, and this metric is analyzed alongside overall accuracy, per-style accuracy, and instruction-following failure rates. All experiments were conducted locally on consumer CPU hardware using unmodified base models without medical fine-tuning, reflecting realistic deployment conditions in low-resource healthcare environments.
The analysis yields several key findings. First, consistency and accuracy emerge as independent performance dimensions: models that produce highly stable answers across prompt variations are not necessarily more accurate. Second, roleplay prompt styles consistently reduce performance across models and datasets, indicating that persona-based prompting may degrade reliability in clinical question answering tasks. Third, instruction-following failures vary independently of model size, suggesting that larger parameter counts do not guarantee more dependable outputs. Finally, Meditron-7B, a domain-pretrained model without instruction tuning, exhibits near-complete instruction-following failure despite encoding substantial medical knowledge, demonstrating that domain knowledge and instruction-following capability are distinct requirements for clinical AI deployment.
The remainder of this paper is structured as follows. Section 2 reviews related work on clinical LLM evaluation, prompt sensitivity, and small language models in healthcare. Section 3 describes the datasets, models, prompt variation design, inference setup, and evaluation metrics. Section 4 presents experimental results. Section 5 discusses clinical implications, limitations, and future directions. Section 6 concludes the paper.
2 Related Work
Most prior work evaluating large language models in healthcare has focused on accuracy, reasoning quality, and output safety. The CLEVER framework (Kocaman and others, 2025) evaluates clinical LLM outputs through blinded physician preference on tasks including summarization and question answering, demonstrating that smaller domain-specific models can outperform larger general-purpose systems on clinically relevant criteria. Kim et al. (Kim and others, 2025) systematically examine medical hallucinations in foundation models, highlighting the dangers of confident but incorrect outputs and their potential to mislead clinical decision-making. While these studies provide critical insights into output quality and safety, neither investigates whether models produce consistent answers when the same clinical question is phrased differently, leaving prompt sensitivity largely unexplored in healthcare AI.
Growing interest in small, CPU-runnable language models has emerged in response to the practical constraints of healthcare deployment in low-resource settings. Garg et al. (Garg and others, 2024) survey small language models in healthcare, documenting their efficiency advantages and suitability for local deployment under limited infrastructure. The MedQA (Jin et al., 2021), MedMCQA (Pal et al., 2022), and PubMedQA (Jin et al., 2019) benchmarks provide standardized evaluation tasks across diverse clinical question formats and have been widely adopted for assessing compact models. These works collectively advance the case for small model deployment in accessible healthcare AI but focus primarily on task accuracy and computational efficiency rather than answer stability under prompt variation.
Prompt sensitivity, defined as the tendency of model outputs to change in response to semantically equivalent but stylistically different inputs, has been studied in general NLP research. Ngweta et al. (Ngweta and others, 2024) demonstrate that minor changes in prompt formatting cause measurable performance fluctuations across multiple models and benchmarks and propose mitigation strategies to reduce this brittleness. However, such investigations have not been extended to clinical question answering, and the safety implications of prompt sensitivity for healthcare AI deployment remain largely unexplored.
Taken together, these lines of research motivate the present study, which systematically evaluates prompt sensitivity and answer consistency in small open-source clinical language models. This work complements prior studies on accuracy, hallucination, and general prompt robustness by focusing on reliability under realistic deployment constraints.
3 Methods
This study follows a controlled empirical evaluation design to assess the reliability of small language models under prompt variation in clinical question answering tasks.
3.1 Datasets
Model consistency and accuracy were evaluated using three established clinical question answering benchmarks. MedQA (Jin et al., 2021) contains multiple-choice questions derived from the United States Medical Licensing Examination (USMLE), representing clinical knowledge required for medical licensure. MedMCQA (Pal et al., 2022) consists of multiple-choice questions from Indian medical entrance examinations (AIIMS/NEET), covering diverse medical specialties including pharmacology, anatomy, and pathology. Although the dataset includes subject labels, the evaluation was conducted on a randomly sampled subset of 200 questions without stratification by specialty. PubMedQA (Jin et al., 2019) presents biomedical research questions paired with PubMed abstracts, requiring yes, no, or maybe answers based on provided evidence rather than memorized knowledge.
From each dataset, 200 questions were randomly sampled using a fixed random seed (seed = 42) to ensure reproducibility. This sample size was selected to balance statistical coverage with the computational constraints of local CPU inference across multiple models and prompt variations, resulting in 15,000 total inference calls.
3.2 Models
Five open-source language models spanning 2B–7B parameters were evaluated, including one domain-adapted medical pretraining model. Table 1 summarises the architectural and training characteristics of each model. The four instruction-tuned models differ in their fine-tuning methodology, pretraining scale, and prompt template conventions in ways that plausibly influence prompt sensitivity behaviour.
Phi-3 Mini (Abdin and others, 2024) (3.8B, Microsoft) is a dense decoder-only transformer fine-tuned using Supervised Fine-Tuning (SFT), Direct Preference Optimisation (DPO), and Reinforcement Learning from Human Feedback (RLHF). Its pretraining relies heavily on high-quality synthetic data and filtered educational content, an approach motivated by the “textbooks are all you need” philosophy. This regime prioritises structured reasoning and instruction adherence; however, the model’s strict chat template format (<|user|>...<|end|>) may create sensitivity to prompts that deviate from this expected structure, contributing to the elevated UNKNOWN rate observed on MedQA.
Llama 3.2 (Grattafiori and others, 2024) (3B, Meta) is derived from Llama 3.1 via pruning and knowledge distillation from larger Llama 3.1 models (8B and 70B), followed by multiple rounds of post-training alignment using SFT, Rejection Sampling (RS), and DPO. This multi-round alignment process, explicitly optimising dialogue helpfulness across varied instruction formats, likely contributes to the model’s strong instruction-following performance and its low UNKNOWN rates across all datasets.
Gemma 2 (Gemma Team and others, 2024) (2B, Google DeepMind) is trained via knowledge distillation from a substantially larger teacher model rather than standard next-token prediction, using on-policy distillation to mitigate train-inference distribution mismatch. Post-training applies SFT, RLHF with a conversational reward model, and model merging across hyperparameter configurations. Knowledge distillation at this scale may instil stable but potentially miscalibrated response tendencies inherited from the teacher model, providing a mechanistic account of Gemma 2’s high consistency scores alongside its low accuracy: the model may be replicating confident response patterns from its teacher that do not generalise correctly to clinical reasoning tasks.
Mistral (Jiang and others, 2023) (7B, Mistral AI) uses sliding window attention and grouped-query attention for efficient inference, and is instruction-tuned using SFT on publicly available conversation datasets. Compared to Llama 3.2 and Phi-3, Mistral’s instruct fine-tuning used a more limited set of public instruction data and requires strict [INST]...[/INST] delimiters. Prompts that do not conform to this format may partially escape instruction alignment, potentially contributing to higher UNKNOWN rates relative to Llama 3.2 and Gemma 2.
Meditron-7B (Chen and others, 2023) (EPFL) is built on Llama-2 and further pretrained on a curated medical corpus comprising PubMed abstracts and international clinical practice guidelines. Critically, Meditron-7B is a continued-pretraining model without instruction tuning; it has not been trained to follow structured task instructions or produce constrained single-token outputs. Its inclusion serves as an architectural control condition: by contrasting a domain-knowledgeable but non-instruction-tuned model against instruction-tuned general models, this study isolates instruction-following capability as a distinct and necessary architectural requirement for structured clinical question answering, independent of domain knowledge.
All models were selected because they are fully open-source, capable of running on consumer CPU hardware without GPU acceleration, and represent distinct model families with meaningfully different training methodologies, enabling cross-architecture comparison of consistency behaviour. All inference was conducted locally using Ollama without domain-specific fine-tuning, ensuring that observed consistency patterns reflect inherent model behaviour rather than task-specific adaptation.
[…]
Model Params Base Model Inst. Tuned Fine-Tuning Method Medical Domain Chat Template
Phi-3 Mini 3.8B Phi-3 (Microsoft) Yes SFT + DPO + RLHF No <|user|>..<|end|>
Llama 3.2 3B Llama 3.1 (Meta) Yes SFT + RS + DPO (multi-round) No <|start_header_id|>..<|eot_id|>
Gemma 2 2B Gemma 2 (Google) Yes KD + SFT + RLHF + Model Merging No <start_of_turn>..<end_of_turn>
Mistral 7B 7B Mistral (Mistral AI) Yes SFT (public datasets) No [INST]..[/INST]
Meditron-$\dagger 7B$ Llama-2 (Meta) No Continued pretraining only (PubMed + guidelines) Yes None
3.3 Prompt Variation Design
To systematically evaluate how small language models respond to differently worded clinical prompts, a prompt variation engine was designed to generate five semantically equivalent but stylistically distinct prompt formulations for each question.
  • Original: The question exactly as it appears in the dataset, serving as the baseline condition.
  • Formal: Rephrased in clinical academic language to simulate how a healthcare professional might query a model in a structured setting.
  • Simplified: Written in plain everyday language, simulating how a non-expert or patient might pose the same question.
  • Roleplay: The model is instructed to respond as a practicing physician, testing whether persona-based prompting affects answer consistency.
  • Direct: Presents only the bare question and answer options with minimal framing, assessing model behavior without explicit instruction.
Each question across all three datasets was transformed into these five prompt styles, yielding 3,000 total prompts (600 questions $\times$ 5 styles). By keeping semantic content identical across variations, any differences in model output can be attributed to prompt sensitivity rather than differences in question content.
3.4 Inference Setup
All models were served locally via Ollama and queried through its REST API. Inference was conducted with temperature set to 0 and a maximum token limit of 10 tokens per response. Setting temperature to 0 ensures deterministic outputs, meaning that identical prompts produce identical responses across runs. This configuration isolates prompt variation as the primary source of output differences rather than stochastic sampling effects. Limiting output to 10 tokens prevents verbose responses and constrains the model to produce the single-letter (A, B, C, or D) or keyword (yes, no, or maybe) answer required by each dataset format.
Model responses were parsed using regular expression extraction. For multiple-choice datasets (MedQA and MedMCQA), the extractor identified the first standalone letter A, B, C, or D in the response. For PubMedQA, the extractor identified the first occurrence of yes, no, or maybe. Responses that did not contain a valid answer were categorized as UNKNOWN, representing instruction-following failures where the model does not produce a valid answer option. All responses, including UNKNOWN outputs, were retained for consistency scoring and subsequent analysis. All inference parameters were held constant across models and prompt styles to ensure fair comparison.
3.5 Evaluation Metrics
Four complementary metrics were defined to quantify model behavior under prompt variation.
Consistency Score. For each question, the majority answer was identified, defined as the response selected most frequently across the five prompt styles, and the proportion of prompt styles that agreed with it was computed. Formally, given responses $r_{1},r_{2},r_{3},r_{4},r_{5}$ for a question, the consistency score was defined as:
$C=\frac{\sum_{i=1}^{5}\mathbf{1}[r_{i}=\hat{r}]}{5}$ (1)
where $\hat{r}$ denotes the majority answer and $\mathbf{1}[\cdot]$ is the indicator function. A score of 1.0 indicates perfect consistency across all prompt styles, while a score of 0.2 indicates maximum inconsistency. For example, responses [A, B, B, B, A] yield a majority answer of B and a consistency score of $3/5=0.60$.
Overall Accuracy. The majority answer for each question is compared with the dataset ground-truth label, reflecting task-level correctness while accounting for variability across prompt styles.
Per-Style Accuracy. Accuracy is computed separately for each prompt style (original, formal, simplified, roleplay, direct), enabling identification of which prompt formulations the model handles most and least reliably.
UNKNOWN Rate. The proportion of responses that do not contain a valid answer option (A–D or yes/no/maybe), representing instruction-following failures that undermine clinical reliability.
Statistical significance was assessed using the Wilcoxon signed-rank test for comparisons of consistency scores and McNemar’s test for paired accuracy comparisons. All p-values are reported without adjustment for multiple comparisons. Because the analysis is exploratory and hypothesis-generating, no correction for multiple testing was applied, consistent with prior exploratory clinical AI benchmarking studies (Bedi and others, 2025). Individual p-values should therefore be interpreted cautiously rather than as confirmatory statistical evidence.
Together, these metrics provide a holistic assessment of model behavior. A model may be highly consistent yet systematically incorrect, accurate only under specific prompt styles, or prone to instruction-following failures. Evaluating any single metric in isolation would obscure these clinically relevant failure modes.
4 Results
4.1 Overall Consistency and Accuracy
[…]
Dataset Model Mean Fully Overall UNKNOWN Acc. Acc. Acc. Acc. Acc.
Cons. Cons. Acc. Rate Orig. Form. Simp. Role. Dir.
MedQA Phi-3 Mini 0.698 19.5 48.0 10.5 44.5 44.5 42.5 26.5 48.0
Llama 3.2 0.776 35.0 49.0 0.8 47.5 44.5 46.0 47.0 43.5
Gemma 2 0.888 63.5 40.0 2.1 39.0 39.0 37.0 35.0 35.0
Mistral 7B 0.800 41.0 45.0 4.7 45.5 41.5 43.0 34.0 50.5
[…]
MedMCQA Phi-3 Mini 0.730 27.0 53.0 4.6 52.0 49.5 48.5 37.5 52.0
Llama 3.2 0.774 38.5 55.5 1.5 56.5 50.5 51.5 48.0 53.0
Gemma 2 0.851 56.5 43.5 0.9 42.5 39.0 38.0 40.5 44.0
Mistral 7B 0.812 45.5 45.5 4.7 44.5 43.0 42.5 41.5 44.0
[…]
PubMedQA Phi-3 Mini 0.830 41.0 48.0 0.0 42.5 53.5 52.5 47.5 31.0
Llama 3.2 0.807 45.5 65.0 2.8 59.5 51.0 64.5 65.0 56.5
Gemma 2 0.845 50.5 33.0 1.3 28.0 41.5 47.5 27.0 30.5
Mistral 7B 0.825 49.5 42.5 7.2 39.5 39.5 45.5 34.5 45.5
[…]
Statistical significance of observed differences was assessed using the Wilcoxon signed-rank test for consistency scores and McNemar’s test for accuracy comparisons. Consistency differences between models were statistically significant on MedQA and MedMCQA (p<<0.001 for most pairwise comparisons) but not on PubMedQA, suggesting that dataset characteristics influence the degree of prompt sensitivity. Accuracy differences were generally non-significant on MedQA, whereas on PubMedQA, Llama 3.2 significantly outperformed both Phi-3 Mini and Gemma 2 (p<<0.001), indicating stronger performance on biomedical evidence-based questions. Full significance test results are provided in Table LABEL:tab:stats.
[…]
Dataset M1 M2 Metric p Sig.
MedQA Llama 3.2 Phi-3 Mini Cons. 0.0005 ***
Gemma 2 Phi-3 Mini Cons. <<.0001 ***
Mistral Phi-3 Mini Cons. <<.0001 ***
Llama 3.2 Gemma 2 Cons. <<.0001 ***
Meditron Phi-3 Mini Cons. <<.0001 ***
Meditron Llama 3.2 Cons. 0.3342 ns
Llama 3.2 Phi-3 Mini Acc. 0.9020 ns
Gemma 2 Phi-3 Mini Acc. 0.0689 ns
Mistral Phi-3 Mini Acc. 0.5383 ns
Llama 3.2 Gemma 2 Acc. 0.0512 ns
Meditron Phi-3 Mini Acc. 0.0004 ***
Meditron Llama 3.2 Acc. 0.0001 ***
MedMCQA Llama 3.2 Phi-3 Mini Cons. 0.0437 *
Gemma 2 Phi-3 Mini Cons. <<.0001 ***
Mistral Phi-3 Mini Cons. 0.0001 ***
Llama 3.2 Gemma 2 Cons. 0.0001 ***
Meditron Phi-3 Mini Cons. 0.0007 ***
Meditron Llama 3.2 Cons. 0.2052 ns
Llama 3.2 Phi-3 Mini Acc. 0.5962 ns
Gemma 2 Phi-3 Mini Acc. 0.0327 *
Mistral Phi-3 Mini Acc. 0.1060 ns
Llama 3.2 Gemma 2 Acc. 0.0075 **
Meditron Phi-3 Mini Acc. 0.0002 ***
Meditron Llama 3.2 Acc. <<.0001 ***
PubMedQA Llama 3.2 Phi-3 Mini Cons. 0.1055 ns
Gemma 2 Phi-3 Mini Cons. 0.8497 ns
Mistral Phi-3 Mini Cons. 0.6172 ns
Llama 3.2 Gemma 2 Cons. 0.1055 ns
[…]
Llama 3.2 Phi-3 Mini Acc. <<.0001 ***
Gemma 2 Phi-3 Mini Acc. 0.0002 ***
Mistral Phi-3 Mini Acc. 0.1930 ns
Llama 3.2 Gemma 2 Acc. <<.0001 ***
Meditron Phi-3 Mini Acc. <<.0001 ***
Meditron Llama 3.2 Acc. <<.0001 ***
Across all three datasets, consistency scores varied across models without a clear correlation with model size (Figure 1). Phi-3 Mini (3.8B) exhibited the lowest mean consistency scores (0.698–0.830), while Gemma 2 (2B) achieved the highest (0.845–0.888) despite being the smallest model tested. Mistral 7B demonstrated strong but slightly variable consistency (0.800–0.825), and Llama 3.2 showed intermediate consistency (0.774–0.807).
Accuracy patterns diverged substantially from consistency rankings (Figure 2). Llama 3.2 achieved the highest overall accuracy across datasets (49.0–65.0%), followed by Phi-3 Mini (48.0–53.0%) and Mistral 7B (42.5–45.5%). Gemma 2, despite exhibiting the highest consistency scores, achieved the lowest accuracy (33.0–43.5%). This inverse relationship between consistency and accuracy in Gemma 2 highlights a critical finding: high consistency does not imply high accuracy. Models can be reliably wrong, producing the same incorrect answer across all prompt variations, which represents a particularly dangerous failure mode in clinical decision support contexts.
Meditron-7B, included as a domain-adapted pretraining model without instruction tuning, exhibited markedly different behavior from all instruction-tuned models. While it achieved moderate consistency on MedQA and MedMCQA (0.761 and 0.762, respectively), its UNKNOWN rates were substantially higher (22.8% and 22.4%), indicating frequent instruction-following failures. On PubMedQA, Meditron-7B produced near-complete failure, with a consistency score of 0.010, 2.0% accuracy, and a 99.0% UNKNOWN rate. This suggests that the model is unable to produce structured yes/no/maybe responses without instruction tuning. These results indicate that domain-specific pretraining alone is insufficient for structured clinical question answering and that instruction-following capability is a prerequisite for clinical deployment.
Due to the extremely high UNKNOWN rate observed for Meditron-7B on PubMedQA (99.0%), statistical comparisons involving this model on that dataset should be interpreted cautiously, as the near-absence of valid responses renders consistency metrics unreliable.
Figure 1: Mean consistency scores across models and datasets. Higher scores indicate greater agreement across prompt styles.
Figure 2: Overall accuracy across models and datasets, calculated using majority answer against ground truth labels.
4.2 Effect of Prompt Style on Accuracy
Prompt style demonstrated a consistent and measurable effect on model accuracy across all five models and three datasets (Figure 3). The Roleplay prompt style, in which models were instructed to respond as a practicing physician, consistently underperformed relative to all other styles. The most pronounced decline was observed in Phi-3 Mini, which achieved 48.0% accuracy under the Direct style but only 26.5% under the Roleplay style on MedQA, representing a drop of 21.5 percentage points. Similar reductions were observed across Llama 3.2, Gemma 2, and Mistral 7B, suggesting that roleplay prompting is systematically detrimental rather than model-specific (Figure 4).
In contrast, the Direct and Original prompt styles produced the highest and most stable accuracy across models and datasets, suggesting that minimal framing is preferable for reliable clinical question answering. These findings carry direct implications for healthcare AI deployment: prompt formulations intended to simulate clinical expertise, such as persona-based instructions, may paradoxically reduce the reliability of small language model outputs in patient-facing applications.
Figure 3: Accuracy by prompt style across models and datasets. Roleplay consistently underperforms relative to other styles.
Figure 4: Roleplay prompt accuracy versus best performing non-roleplay style across all models and datasets, with annotated performance gaps.
4.3 Instruction-Following Failure Rate
The UNKNOWN rate, defined as the proportion of responses that did not contain a valid answer option, varied substantially across models and datasets (Figure 5). Phi-3 Mini exhibited the highest UNKNOWN rate on MedQA at 10.5%, indicating that approximately one in ten queries failed to produce a usable response. Mistral 7B, despite being the largest model evaluated, also demonstrated non-negligible UNKNOWN rates across all three datasets (4.7%–7.2%), suggesting that model size alone does not guarantee reliable instruction following.
In contrast, Llama 3.2 and Gemma 2 achieved the lowest UNKNOWN rates across datasets (0.8%–2.8% and 0.9%–2.1%, respectively), consistently producing responses in the expected format. These results suggest that instruction adherence is a model-specific characteristic independent of parameter count.
Meditron-7B exhibited the most severe instruction-following failures of all models evaluated, with UNKNOWN rates of 22.8% on MedQA, 22.4% on MedMCQA, and 99.0% on PubMedQA. The near-complete failure on PubMedQA, where 99 of 100 responses were invalid, is consistent with Meditron’s architecture as a base pretraining model without instruction tuning. Unlike the instruction-following failures observed in Phi-3 Mini, which appear to be prompt-style dependent, Meditron’s failures reflect a fundamental absence of structured response capability rather than prompt sensitivity.
From a clinical deployment perspective, high UNKNOWN rates represent a critical reliability concern. A model that frequently fails to produce a valid response cannot serve as a dependable decision support tool, as such failures interrupt clinical workflows, introduce uncertainty, and erode clinician trust in AI-assisted systems.
The variation in UNKNOWN rates across instruction-tuned models also reflects differences in the scope and methodology of their post-training alignment. Llama 3.2 and Gemma 2, which produced the lowest UNKNOWN rates, underwent multi-round alignment processes optimising across diverse instruction formats. Mistral 7B and Phi-3 Mini, which showed higher UNKNOWN rates on certain datasets, relied on more constrained instruction fine-tuning that may be more sensitive to prompt phrasings that diverge from their expected template structures. These patterns suggest that the breadth of instruction format coverage during fine-tuning, not merely parameter count, is the primary architectural determinant of instruction-following reliability in structured clinical tasks.
Figure 5: Instruction-following failure rates (UNKNOWN responses) across models and datasets. Phi-3 Mini shows the highest failure rate on MedQA at 10.5%.
4.4 Consistency Versus Accuracy Relationship
Analysis of the relationship between mean consistency scores and overall accuracy revealed no clear positive correlation across models and datasets (Figure 6). Gemma 2 exemplifies this disconnect most clearly: it achieved the highest consistency scores across all datasets (0.845–0.888) while simultaneously recording the lowest accuracy (33.0–43.5%), indicating that it consistently produces incorrect answers with high confidence. Conversely, Llama 3.2 demonstrated moderate consistency (0.774–0.807) yet achieved the highest accuracy across most datasets (49.0–65.0%), showing that a model can be more frequently correct even when its answers vary across prompt styles.
The distribution of consistency scores further illustrates model-level differences in reliability (Figure 7). Phi-3 Mini exhibited the widest spread of consistency scores on MedQA, indicating substantial variability in prompt sensitivity across questions. Gemma 2 showed a tighter, higher distribution, remaining consistently stable but systematically inaccurate.
These findings demonstrate that consistency and accuracy are independent dimensions of model performance. Evaluating either metric in isolation is insufficient for clinical AI assessment. A model that is highly consistent but consistently wrong may be more dangerous than one that occasionally contradicts itself, as the former provides false confidence in systematically incorrect outputs. Safe clinical deployment therefore requires joint evaluation of both consistency and accuracy.
Figure 6: Scatter plot of mean consistency score versus overall accuracy across all models and datasets. No clear positive correlation is observed, indicating that consistency and accuracy are independent metrics.
Figure 7: Distribution of consistency scores across models and datasets. Phi-3 Mini shows the widest spread on MedQA; Gemma 2 shows a tight but high distribution reflecting systematic consistency without accuracy.
5 Discussion
From a clinical informatics perspective, these findings highlight the importance of evaluating AI systems not only for accuracy but also for response stability. Clinical decision support tools are often queried in multiple ways by clinicians, and inconsistent responses to semantically equivalent queries may undermine trust in AI-assisted systems. Conversely, consistently incorrect outputs may create a false sense of reliability if repeated answers are interpreted as evidence of model confidence. Evaluation frameworks that jointly assess accuracy and response stability therefore provide a more comprehensive basis for assessing whether language models are sufficiently reliable for integration into clinical workflows.
5.1 Consistency and Accuracy are Independent Metrics
A central empirical finding of this study is that consistency and accuracy represent independent dimensions of small language model performance in clinical question answering. While prior evaluations of clinical AI have predominantly focused on accuracy as the primary metric (Bedi and others, 2025), the results presented here demonstrate that a model can exhibit high consistency while still producing systematically incorrect answers, a failure mode referred to in this work as reliable incorrectness.
Gemma 2 illustrates this most clearly, achieving the highest mean consistency scores across all datasets (0.845–0.888) while recording the lowest overall accuracy (33.0–43.5%). This pattern suggests that the model has learned stable response tendencies that are nonetheless clinically incorrect. Across 200 MedQA questions, Gemma 2 produced perfectly consistent yet incorrect answers on 77 questions (38.5%), selecting the same wrong option across all five prompt styles with a consistency score of 1.0.
The danger of reliable incorrectness is both psychological and technical. Consider a primary care physician using an AI assistant to support diagnostic reasoning. If the model returns the same answer regardless of how a question is phrased, this repetition creates an illusion of reliability. A clinician might reason, “it keeps suggesting pulmonary embolism; it must be confident.” That apparent stability can feel reassuring. However, if the correct diagnosis is pneumonia, the model’s consistency actively reinforces a false conclusion. Over time, clinicians may develop misplaced trust in systems that appear stable and decisive without recognising that the stability reflects systematic error rather than clinical knowledge. In practice, this could translate into repeated misdiagnoses, inappropriate investigations, or delayed treatment, not because the model behaves randomly, but because it is confidently and predictably incorrect.
This finding aligns with Kim et al. (Kim and others, 2025), who demonstrate that foundation models frequently produce hallucinations with high apparent confidence and that clinicians consider such outputs particularly dangerous because they are difficult to detect. The present results extend this concern to the domain of prompt sensitivity. A model that produces consistent but incorrect answers across varied prompt formulations compounds this risk by appearing robust to rephrasing.
These observations underscore that accuracy and consistency must be evaluated jointly in any clinical AI assessment framework. Neither metric alone is sufficient to characterize the safety profile of a model intended for healthcare deployment.
The architectural characteristics of Gemma 2 offer a plausible mechanistic account of this pattern. The 2B Gemma 2 model is trained via knowledge distillation from a substantially larger teacher model, inheriting the teacher’s response tendencies rather than learning independently from ground-truth labels. If the teacher model has developed systematic biases in clinical question answering — for instance, preferring certain option positions or applying heuristics that generalise poorly to medical reasoning — these biases may be faithfully replicated in the distilled student model, producing high cross-prompt consistency that reflects the teacher’s error patterns rather than genuine clinical knowledge. This distillation-induced miscalibration represents a distinct architectural risk factor for clinical AI deployment that accuracy-only evaluation frameworks would not detect.
5.2 Roleplay Prompts Consistently Underperform
Across all five models and three datasets, roleplay-style prompts consistently produced lower accuracy than direct or original question formats. The most pronounced decline was observed in Phi-3 Mini on MedQA, where roleplay accuracy dropped 21.5 percentage points below the best-performing style. Critically, this pattern was consistent across all model families tested, suggesting a systematic rather than model-specific phenomenon.
Roleplay prompts appear to introduce a form of task interference that is particularly detrimental for small language models. When a prompt begins with persona framing such as “You are a senior physician taking a licensing examination,” the model must simultaneously interpret and simulate a clinical identity while reasoning through a structured medical question.
For smaller models with limited representational capacity, this dual requirement may dilute attention from the core reasoning task. Furthermore, persona-based prompts may activate patterns learned from conversational or narrative training data rather than the structured exam-style reasoning required for clinical QA benchmarks, effectively shifting the task from question answering to performance-oriented text generation.
This finding has direct practical implications. Prompt engineering guidelines for clinical AI systems often recommend persona framing to make outputs sound more authoritative or clinically appropriate (White and others, 2023). These results suggest that, for small open-source models, such framing may paradoxically reduce factual reliability. Developers building clinical decision support tools on small language models should favor minimal, direct prompt formulations over persona-based instructions.
This observation also extends findings from Ngweta et al. (Ngweta and others, 2024), who demonstrated that prompt format changes cause measurable performance variation in general NLP tasks. The present results show that, in clinical settings, one specific prompt format, roleplay, is systematically and substantially more harmful than others, representing a concrete and actionable finding for healthcare AI deployment.
This phenomenon is likely amplified by the prompt template conventions enforced during instruction fine-tuning. Each model evaluated uses a distinct chat template: Phi-3 Mini employs <|user|>...<|end|> delimiters, Llama 3.2 uses <|start_header_id|>...<|eot_id|> markers, and Mistral 7B requires [INST]...[/INST] framing. Roleplay prompts, which begin with persona instructions such as “You are a senior physician,” may partially misalign with the expected input structure of these templates, shifting the model’s processing toward narrative or conversational generation rather than constrained structured reasoning. This suggests that roleplay sensitivity reflects not merely task framing but an interaction between prompt style and the learned template expectations embedded during instruction tuning — an architectural rather than purely behavioural phenomenon.
5.3 Implications for Low-Resource Healthcare Deployment
A primary motivation for this study was the practical challenge of deploying AI in resource-constrained healthcare environments. Many rural clinics, community hospitals, and health systems in low-income settings lack access to GPU infrastructure, reliable high-speed internet, or cloud-based AI services. In these contexts, locally deployable open-source models in the 2B–7B parameter range represent a more realistic and equitable alternative to large proprietary systems. The experimental design intentionally reflected these constraints: all models were evaluated on consumer CPU hardware without domain-specific fine-tuning, simulating realistic deployment conditions rather than idealized research environments.
These findings provide concrete guidance for model selection in such settings. Although Gemma 2 demonstrated the highest consistency scores, its substantially lower accuracy suggests that it may systematically mislead clinicians, particularly in low-oversight environments where AI outputs may not be routinely verified. Mistral 7B, despite being the largest model evaluated, did not clearly outperform smaller alternatives and exhibited unexpected instruction-following failures, suggesting that larger parameter counts do not automatically justify greater hardware demands.
Llama 3.2 (3B) emerged as the most balanced candidate under the evaluation criteria used in this study, achieving the highest overall accuracy (49.0–65.0%), moderate consistency, and among the lowest instruction-following failure rates. For low-resource healthcare deployment scenarios where clinician oversight may be limited, a model offering strong accuracy and reliable instruction adherence is preferable to one that maximizes consistency at the expense of correctness.
More broadly, these results suggest that deployment decisions for clinical AI systems should be guided by joint evaluation of accuracy, consistency, and instruction compliance rather than any single metric. In settings where errors carry direct patient safety implications, the cost of systematic incorrectness, even when delivered consistently, outweighs the benefit of apparent stability. This study provides a practical evaluation framework for low-resource healthcare AI deployment that can be reproduced without specialized infrastructure, making it accessible to researchers and practitioners in the settings it is intended to serve.
5.4 Domain Knowledge Versus Instruction Following
The inclusion of Meditron-7B reveals an important distinction between two capabilities required for clinical AI deployment: domain knowledge and instruction adherence. Meditron-7B was pretrained on a curated medical corpus comprising PubMed articles and clinical guidelines (Chen and others, 2023), encoding substantial medical domain knowledge. However, without instruction tuning, it failed to produce structured responses in the required format across all three datasets, with UNKNOWN rates reaching 99.0% on PubMedQA.
This finding suggests that domain knowledge and task usability are orthogonal capabilities. A model can encode rich medical knowledge through continued pretraining yet remain clinically unusable for structured question answering if it lacks the ability to follow task-specific instructions. For practitioners deploying AI in clinical settings, this distinction has direct practical implications. Selecting a model based solely on its training data, without verifying instruction-following capability, may result in a system that is knowledgeable but functionally unusable in structured workflows.
These observations also motivate a clear direction for future work: instruction-tuning Meditron or similar domain-pretrained models on structured clinical QA datasets and evaluating whether the combination of domain knowledge and instruction following produces models that are both more accurate and more consistent than general instruction-tuned alternatives.
From an architectural standpoint, this finding reflects a fundamental distinction between two training objectives that are often conflated in clinical AI discourse. Continued pretraining on medical corpora, as applied to Meditron-7B, optimises the model’s internal representations of medical knowledge but does not instil the structured output behaviour required for task-constrained deployment. Instruction tuning, by contrast, explicitly trains the model to produce constrained responses in specific formats in response to specific prompt structures. These are orthogonal capabilities: a model can encode rich biomedical knowledge through pretraining yet remain clinically unusable because it lacks the architectural conditioning to translate that knowledge into a structured, deployable output. For practitioners selecting models for low-resource clinical deployment, this distinction implies that medical domain pretraining alone is an insufficient selection criterion; instruction-following capability must be verified independently.
5.5 Limitations
Several limitations should be considered when interpreting these findings. These limitations define the scope of the current study and collectively motivate the future directions described in Section 5.6.
First, 200 questions per dataset were sampled, which provides a computationally feasible and reproducible evaluation but may not fully capture the diversity of each benchmark. Second, the three datasets used, MedQA, MedMCQA, and PubMedQA, focus primarily on structured examination-style and research-based questions and may not reflect the complexity of real-world clinical dialogue or unstructured patient interactions.
Third, human clinical evaluation was not conducted to assess whether consistent but incorrect model outputs would meaningfully influence physician decision-making in practice. Such evaluation would strengthen the clinical validity of these findings but was beyond the scope of this study. Fourth, all models were evaluated without domain-specific fine-tuning, which may underestimate the performance achievable through medical adaptation techniques such as instruction tuning or retrieval-augmented generation (Lewis and others, 2020).
Fifth, the evaluation was restricted to multiple-choice and yes/no/maybe answer formats, whereas real clinical environments frequently involve open-ended reasoning, differential diagnosis, and contextual patient data. Sixth, hardware constraints limited the evaluation to models in the 2B–7B parameter range running on consumer CPU hardware, precluding direct comparison with larger proprietary systems such as GPT-4 or Claude. Seventh, while the five prompt styles were designed to simulate realistic variation in clinical query formulation, they do not exhaust the full range of phrasings a clinician might use in practice. Finally, Meditron-7B was evaluated in its base pretraining form without instruction tuning, which limits direct comparison with instruction-tuned models and may underrepresent the potential of domain-adapted models when properly instruction-tuned for structured clinical tasks.
Together, these limitations define the scope of this study as an initial systematic analysis of prompt sensitivity under constrained deployment conditions. The evaluation framework introduced here is designed to be reproducible and extensible, supporting future clinical validation studies without requiring specialised infrastructure.
5.6 Future Work
Several directions emerge naturally from this study. First, future research should evaluate whether domain-specific fine-tuning of small models on medical corpora reduces prompt sensitivity and improves consistency. If fine-tuned models demonstrate stronger stability under prompt variation, this would suggest that training data alignment, rather than model architecture, is the primary driver of inconsistency.
In particular, instruction-tuning Meditron-7B or similar domain-pretrained models on structured clinical QA datasets represents a direct next step. This would test whether combining domain knowledge with instruction-following capability produces models that are both more accurate and more consistent than the general instruction-tuned alternatives evaluated in this study.
Second, integrating retrieval-augmented generation (RAG) pipelines with small clinical models represents a promising direction. RAG-augmented systems ground responses in retrieved evidence, which may reduce both factual hallucinations and prompt-induced answer variation. Evaluating whether external knowledge retrieval mitigates the consistency–accuracy tradeoff observed in this study would be a natural extension (Lewis and others, 2020).
Third, evaluating larger open-source and proprietary models, when hardware resources permit, would clarify whether the independence between consistency and accuracy observed here persists at greater parameter scales or whether it is specific to sub-7B models.
Beyond technical extensions, human-centered validation is essential. A controlled clinical study involving practicing physicians could assess how consistent but incorrect model outputs influence clinical reasoning, trust calibration, and decision-making in realistic workflows. Such a study would strengthen the ecological validity of the consistency metric introduced here.
Translationally, these findings support the development of a lightweight clinical decision support prototype for general practitioners in low-resource settings, built around the model demonstrating the strongest balance of accuracy and instruction adherence, with prompt engineering guardrails informed by the sensitivity analysis. Extending this evaluation framework to multilingual datasets and non-English clinical benchmarks is also critical, as many low-resource healthcare environments operate outside English-dominant settings. Together, these directions would advance this work from controlled benchmarking toward safe, deployable clinical AI systems grounded in both technical rigor and real-world constraints.
6 Conclusion
This study systematically evaluates prompt sensitivity in five open-source language models (2B–7B parameters), including four general instruction-tuned models and one domain-adapted pretraining model, across three clinical question answering datasets using five controlled prompt style variations, yielding 3,000 total evaluation instances.
Four principal findings emerge from this analysis. First, consistency and accuracy are independent metrics: models that achieve high consistency are not necessarily more accurate, with Gemma 2 demonstrating the most pronounced case of reliable incorrectness, achieving the highest consistency scores while recording the lowest accuracy across all datasets. Second, roleplay-style prompts consistently reduce accuracy across all models and datasets, suggesting that persona-based framing introduces systematic instability in clinical reasoning tasks and should be avoided in healthcare AI prompt design. Third, instruction-following failures vary substantially across models and are not determined by parameter count alone, indicating that model size is an insufficient proxy for deployment reliability under hardware constraints. Fourth, Meditron-7B, a domain-adapted model without instruction tuning, exhibited near-complete failure on structured clinical QA tasks despite encoding substantial medical domain knowledge, demonstrating that domain knowledge and instruction-following capability are orthogonal requirements for clinical AI deployment.
Taken together, these results demonstrate that consistency alone cannot serve as a proxy for clinical reliability. Models that appear stable under prompt variation may still pose significant patient safety risks if their answers are systematically incorrect. Safe deployment of AI in clinical settings, particularly in low-resource environments where human oversight may be limited, requires joint evaluation of accuracy, consistency, and instruction adherence.
More broadly, this work highlights the need for the research community to move beyond accuracy-only evaluation frameworks for clinical AI and adopt multidimensional assessment approaches that reflect the reliability requirements of real-world healthcare deployment. The evaluation framework, datasets, and code introduced in this study are publicly available at (Hariprasad, 2026) to support reproducibility and future research in this direction.
Acknowledgments
This study received no funding. All experiments were conducted on personal hardware by an independent researcher. The author thanks the open-source communities behind Ollama, HuggingFace Datasets, and the model developers whose publicly released technical reports informed the architectural analysis in this work.
Conflicts of Interest
The author declares no conflicts of interest. No funding was received for this study.
Data Availability
The evaluation framework, scored datasets, and all code used in this study are publicly available at https://github.com/shravani-01/clinical-llm-eval (Hariprasad, 2026). The three clinical benchmarks used (MedQA, MedMCQA, PubMedQA) are publicly available via HuggingFace Datasets.
Ethics Statement
This study used publicly available, de-identified benchmark datasets (MedQA, MedMCQA, PubMedQA) and did not involve human subjects, patient data, or any form of primary data collection. No ethics board review was required.
Appendix A Representative Prompt Variation Example
This appendix presents a single representative example illustrating how the five prompt styles were applied to an identical clinical question, along with the raw model responses and extracted answers from Phi-3 Mini. This example was selected because it demonstrates prompt sensitivity in practice: the model produces the correct answer under three prompt styles (Formal, Simplified, Direct) but selects an incorrect answer under both the Original and Roleplay styles, despite identical question content across all five variations.
Question
A healthy 23-year-old male is undergoing an exercise stress test as part of his physiology class. If blood were to be sampled at different locations before and after the stress test, which area of the body would contain the lowest oxygen content at both time points?
Options:
  • A:
    Inferior vena cava
  • B:
    Coronary sinus (correct answer)
  • C:
    Pulmonary artery
  • D:
    Pulmonary vein
Prompt Texts and Model Responses (Phi-3 Mini)
Table 4: Prompt texts, raw model output, and extracted answers for Phi-3 Mini on a representative MedQA question. The correct answer is B (Coronary sinus). The Roleplay style produces an incorrect answer (A) despite identical question content across all five variations.
Style Prompt Text (abbreviated) Raw Response Extracted Correct?
Original “Answer the following medical question by choosing the correct option. …Answer with only the option letter (A, B, C, or D).” Option A: Inferior vena c… A No
Formal “You are a medical expert. Based on your clinical knowledge, select the most appropriate answer …Respond with only the letter of the correct choice (A, B, C, or D).” B: Coronary sinus B Yes
Simplified “Read this medical question carefully and pick the best answer. …Which letter is correct? Reply with just A, B, C, or D.” B: Coronary sinus The… B Yes
Roleplay “You are a physician taking a medical licensing exam. Answer this question as you would on the exam. …Your answer (A, B, C, or D):” The correct answer is A: Inferior v… A No
Direct “Medical question: …What is the correct option? State only the letter.” B B Yes
The majority answer across five styles is B (3 of 5 responses), yielding a consistency score of $3/5=0.60$ and a correct majority answer. However, the Roleplay prompt produces answer A in both instances it appears, illustrating how persona-based framing systematically shifts the model toward an incorrect response. The Original prompt also returns A, despite minimal structural difference from the Direct prompt — a further illustration of prompt sensitivity even between superficially similar formulations. The raw responses also reflect a secondary finding: the 10-token output limit causes truncation (Option A: Inferior vena c…), contributing to the UNKNOWN rate observed in cases where truncation prevents a valid letter from being extracted.
Full prompt texts for all five styles across all three datasets are available in the project repository [Hariprasad, 2026].
Example 2: Reliable Incorrectness (Gemma 2, Consistency = 1.0, Accuracy = 0)
This example illustrates the reliable incorrectness failure mode described in Section 5.1. Gemma 2 selects the same incorrect answer (D) across all five prompt styles, yielding a perfect consistency score of 1.0 while being entirely wrong. The correct answer is B (Isosorbide dinitrate), consistent with the presentation of Prinzmetal angina, which is treated with nitrates and calcium channel blockers rather than beta-blockers such as propranolol (D). This case demonstrates that a model appearing maximally stable under prompt variation can simultaneously be maximally unreliable from a clinical standpoint.
Question
A 38-year-old woman presents to her physician’s clinic for recurrent episodes of chest pain that wakes her from her sleep. While usually occurring late at night, she has also had similar pains during the day at random times, most recently while sitting at her desk in her office and at other times while doing the dishes at home. The pain lasts 10–15 minutes and resolves spontaneously. She is unable to identify any common preceding event to pain onset. The remainder of her history is unremarkable and she takes no regular medications. Examination reveals: pulse 70/min, respirations 16/min, and blood pressure 120/70 mmHg. A physical examination is unremarkable. Which of the following would be effective in reducing her symptoms?
Options:
  • A:
    Aspirin
  • B:
    Isosorbide dinitrate (correct answer)
  • C:
    Heparin
  • D:
    Propranolol
Model Responses (Gemma 2)
Table 5: Gemma 2 responses across all five prompt styles for a representative MedQA question demonstrating reliable incorrectness. The correct answer is B (Isosorbide dinitrate). Gemma 2 selects D (Propranolol) under every prompt style, yielding a consistency score of 1.0 and 0% accuracy on this question.
Style Prompt Text (abbreviated) Raw Response Extracted Correct?
Original “Answer the following medical question by choosing the correct option. …Answer with only the option letter (A, B, C, or D).” D D No
Formal “You are a medical expert. Based on your clinical knowledge, select the most appropriate answer …Respond with only the letter of the correct choice (A, B, C, or D).” D D No
Simplified “Read this medical question carefully and pick the best answer. …Which letter is correct? Reply with just A, B, C, or D.” D D No
Roleplay “You are a physician taking a medical licensing exam. Answer this question as you would on the exam. …Your answer (A, B, C, or D):” The correct answer is **(D) Propran… D No
Direct “Medical question: …What is the correct option? State only the letter.” D D No
Consistency score = $5/5=1.0$; majority answer = D (incorrect). This pattern — perfect stability, zero correctness — exemplifies why consistency alone is insufficient as a clinical safety metric. A clinician or automated system relying on answer stability as a proxy for confidence would receive maximally misleading signals from this model on this question. Gemma 2 produced 77 such cases on MedQA alone (out of 200 questions), representing 38.5% of its evaluation set.
Full prompt texts for all five styles across all three datasets are available in the project repository [Hariprasad, 2026].